OpenStudy (anonymous):

simplify sqrt75-4sqrt-3sqrt32

OpenStudy (anonymous):

\[\sqrt{75}- 4\sqrt{8}-3\sqrt{32}\]

OpenStudy (anonymous):

\[5\sqrt{3} - 20\sqrt{2}\]